OpenStudy (anonymous):

cos² θ + cos²(π/2 - θ) = 1 prove the identity

OpenStudy (anonymous):

can anybody do it?

OpenStudy (anonymous):

cos^2(pi/2 -t ) = [cos(pi/2 -t)]^2 = [cos(pi/2)cos(t) + sin(pi/2)sin(t)]^2 = sin^2(t) and cos^2(t) + sin^2(t) = 1 done

OpenStudy (anonymous):

good job. Thanks